In the study of 10 patients treated for a kidney infection, the following data refer to the use of urinary analysis 1, 2, 5, 7, 3, 1, 4, 3, 6, and 4 Use the given data to conduct a 99% confidence interval for the mean use of this procedure.
Answer #1

For the given data sample mean is \overline{x}=\frac{\sum x}{n}=3.6

Create the following table.

data data-mean (data - mean)2
1 -2.6 6.76
2 -1.6 2.56
5 1.4 1.96
7 3.4 11.56
3 -0.6 0.36
1 -2.6 6.76
4 0.4 0.16
3 -0.6 0.36
6 2.4 5.76
4 0.4 0.16

Find the sum of numbers in the last column to get. \sum (xi-\overline{x})^2=36.4

So standard deviation is \sigma=\sqrt{\frac{\sum (xi-\overline{x})^2}{n-1}}=2.0111

Now t table value for 9 df is 2.262

So Margin of Error is E=t*\frac{s}{\sqrt{n}}= 3.250*\frac{2.0111}{\sqrt{10}}=2.067

Hence CI is CI=\overline{x} \pm E=3.6 \pm 2.067=(1.533,5.667)